# Clinicians’ Voice: Fundamental Considerations for XAI in Healthcare

### v1

@@ Abstract
Explainable AI (XAI) holds the promise of advancing the implementation and adoption of AI-based tools in practice, especially in high-stakes environments like healthcare. However, most of the current research is disconnected from its practical applications and lacks input of end users. To address this, we conducted semi-structured interviews with clinicians to discuss their thoughts, hopes, and concerns. We find that clinicians generally think positively about developing AI-based tools for clinical practice, but they have concerns about how these will fit into their workflow and how it will impact clinician-patient relations. We further identify education of clinicians on AI as a crucial factor for the success of AI in healthcare and highlight aspects clinicians are looking for in (X)AI-based tools. In contrast to other studies, we take on a holistic and exploratory perspective to identify general requirements, which is necessary before moving on to testing specific (X)AI products for healthcare.
Keywords explainable AI, trustworthy AI, human-centered explainable AI, healthcare, interview study
@@ 1 INTRODUCTION
Rapid developments in artificial intelligence (AI) call for more transparency of AI-based technologies to tackle concerns related to trustworthiness, responsibility, and accountability. This has led to the emergence of the field of Explainable AI (XAI). The goal of XAI is making AI models more transparent and understandable to address these concerns and aid the implementation of AI-based technologies. Especially in high-stakes situations, like those in healthcare, AI-based systems could enhance critical decision-making processes. Ideally, AI models provide a certain prediction (e.g., the probability of a skin lesion being cancerous) which supports the clinicians in their decision process. However, beyond regulatory requirements these technologies have to meet, there is an apparent reluctance to use AI-based tools as it is unclear how the AI model arrives at its prediction. Although user studies, evaluations, and multidisciplinary research are generally advocated for within XAI, only a small fraction of publications connect their work with practical applications. With this research, we contribute to the research on XAI in healthcare by conducting semi-structured interviews with clinicians from the Netherlands. These interviews are aimed at developing a better understanding of the current state of (X)AI in healthcare, and clinicians’ perception of the development of AI-based tools. We identify clinicians’ hopes and concerns for (X)AI in healthcare and formulate concrete takeaways to inform research and development in the field. Aspects we looked at include desirable characteristics of XAI methods/explanations, tasks that clinicians (would) find (X)AI helpful for, concerns and fears they face regarding the implementation of (X)AI in their workflow, and their perception of the role of trust in the implementation and adoption of (X)AI. We confirm findings from previous research, including that users emphasize the importance of receiving information about the background of the model in addition to its predictions and explanations. Our results further highlight the importance of multidisciplinary projects to ensure seamless introduction of (X)AI-based tools into clinical practice, and the significance of educating clinicians on AI before implementing AI-based tools in their workflow.
@@ 2 BACKGROUND
One can generally distinguish between glass-box and black-box models in XAI. Glass-box models are transparent, and model internals and operations are accessible to the users such that they can trace back every prediction made by the model. Thus, glass-box models provide both local and global explainability. Black-box models are opaque machine learning models where the user is unable to trace back the prediction due to the model complexity. Typically, (only) local explainability is achieved by complementing single predictions made by black-box models with post-hoc explanations. For an in-depth account of developments in XAI, we refer to several recent review papers (see, for example, Das et al., 2020; Gilpin et al., 2018; Murdoch et al., 2019; Ras et al., 2022; Linardatos et al., 2021; Guidotti et al., 2018; Carvalho et al., 2019; Barredo Arrieta et al., 2020; Ali et al., 2023)
@@
In healthcare, critical decisions are made daily and hence, the potential for AI-based decision support systems is undeniable. This is also reflected in the rising number of publications on AI-based clinical decision support tools (DSTs) that incorporate explainability. For example, Thoral et al. (2021) trained an AI-model to predict re-admission or death within seven days after discharge from the ICU and explain predictions using feature importance scores Lundberg and Lee (2017). Both Abdullah et al. (2021) and Antoniadi et al. (2021) provide reviews of explainable machine learning approaches in healthcare including examples of publications on medical AI tools with explainability. Markus et al. (2021) focus on how explainability can be evaluated and provide a step-by-step guide to choose the most appropriate XAI method. Amann et al. (2020) take a multidisciplinary approach to determining the need for explainability in healthcare from technological, legal, medical, patient, and ethical perspectives. They conclude that there is an apparent need for explainibility and amongst others highlight the importance of considering the interaction between healthcare professionals and AI-based DSTs.
@@
Cai et al. (2019) conduct a user study including semi-structured interviews on AI-assisted diagnosis of prostate cancer, focusing on identifying desired aspects for explainability in the AI tool. They find that pathologists are interested in global properties of the AI tool beyond accuracy, such as how conservative or liberal its decisions are or what kind of data it was trained on. In a similar fashion, Naiseh et al. (2023) empirically evaluated the effect of different types of explanations on trust calibration in a medical context and conducted interviews with some participants after the experiment. Like Cai et al. (2019), they find that practitioners are interested in global information of AI-tools, such as the contents of the data and the recency of the trained model. They further find that practitioners worry about fitting AI-tools into their workflow because it adds extra workload and that they would like more guidance on how to interpret the explanations. Tonekaboni et al. (2019) conducted interviews with 10 clinicians that are familiar with AI-based clinical tools, primarily to identify what explanation types best serve their needs and to propose evaluation metrics of explainability tools. For example, they found that clinicians value feature importance scores as they draw attention to specific patient characteristics relevant for the prediction. They further conclude that example-based explanations are less helpful as they depend on the definition of similarity with the added difficulty that seemingly similar patients may have significantly different outcomes and vice versa. They also find that accuracy alone is not sufficient and that the output should be complemented with a certainty score for the prediction.
@@
Other studies in the realm of XAI in healthcare have a slightly different focus. For example, Lee et al. (2024) address onboarding material for XAI-tools, Subramanian et al. (2024) focus on a broader stakeholder perspective not only involving clinicians, and Bienefeld et al. (2023) analyse collaboration between developers and clinicians. Our study differs from this previous research in several ways. First, we do not build our work around a specific case or AI-tool. In doing so, we eliminate the factor of expertise from the study design. Second, we interview clinicians from different fields within healthcare to record general data. Third, our study is exploratory in nature rather than empirically driven. That is, we do not test for an effect of XAI methods on, for example, trust or understanding. The goal of our research is to identify clinicians’ requirements for successful implementation of AI-tools in clinical practice with the aim of improving decision making in healthcare.
@@ 3 METHODOLOGY
We conducted 11 semi-structured interviews with open-ended questions with clinicians from the Netherlands practicisng in various fields. The study was approved by the ethical body of our institution and all participants gave their informed consent prior to participation. An interview scheme with prompts and guiding questions was used as basis for the interviews (see Table 1). All interviews were conducted by the same interviewer to limit threats to validity and reliability introduced by interviewer variability. Upon completion of the interviews, each interview was transcribed and then analysed using a coding scheme. We did not pre-define any coding categories as this is an exploratory study. Instead, we collected relevant quotes from the transcribed interviews for each of the topics addressed and then identified common themes. A more detailed account of our methodology including participants and recruitment, procedure, materials, analysis, and limitations is given in Appendix A.
@@ 4.1 Current use of AI in healthcare
Based on the interviews, we can group AI-based tools into those related to (1) business operations, (2) care processes, and (3) medical device regulations. The first category of AI systems is relatively universal to any larger business, not only hospitals or care facilities, and covers aspects like finance, forecasting, and so on. The second category relates to AI systems that enhance the care process by, for instance, making it more efficient. A typical example that several interviewees provided was the planning of chemotherapy. The last category of AI systems in healthcare concerns AI systems that fall under the medical device regulation. Simply put, a medical device is any instrument intended to be used to diagnose, treat, monitor or prevent diseases “and does not achieve its primary intended action by pharmacological, immunological or metabolic means” (International Medical Device Regulators Forum (2013), IMDRF, p.7). AI models fall under the category Software as a Medical Device, which includes software that is intended to be used for any of the aforementioned purposes without being part of a hardware medical device (International Medical Device Regulators Forum (2013), IMDRF). In layman’s terms, decision support tool is more commonly used to describe AI as a medical device. Here, an AI model is trained to make a prediction based on patient data and hence, supports clinicians in their decisions.
When asked to what extent AI models are implemented in medical care in their field specifically, interviewees unanimously said that there is very little AI (as DSTs) integrated in their workflows. Most interviewees pointed out that a significant amount of budget is invested in research and development of medical AI models. Though many of these efforts are being tested, very few of them actually reach the bedside. Arguably, the most popular field for AI applications within healthcare is radiology, with the number of publications on AI in radiology being “sky high” (interview 4). However, even though there are more than 200 products in AI for radiology available on the European market [^1] https://radiology.healthairegister.com, not many of those are integrated in clinical practice yet. Furthermore, one interviewee pointed out that most of these products perform specific subtasks, such as segmentation or quantification tasks, instead of predicting a concrete medical outcome; thus, only indirectly supporting clinicians.
@@ 4.2 What tasks should AI be used for?
Regarding AI as a clinical DST, four interviewees mentioned predicting complications as a useful application. For example, when doing surgery, they would like to know upfront how big the chance is to run into a certain complication. And even beyond that, one interviewee said they “would like to know upfront what can go wrong and how can we prevent that” (interview 5). Hence, AI could serve not only as a warning system but also as a guide on how to navigate certain situations. To give a concrete example, one interviewee from obstetrics and gynaecology explained that when a woman is in labor, they monitor both the mother and the baby, meaning they collect a large dataset. Having an AI model trained on this data that continuously predicts the chance of a birth ending in a C-section would save them not only time and resources, but would also reduce the risks for women in labor. Another example comes from oncology, where it would be beneficial to predict the likelihood of complications during surgery but also the probability of remission following treatment.

### v2

@@ Abstract
Explainable AI (XAI) holds the promise of advancing the implementation and adoption of AI-based tools in practice, especially in high-stakes environments like healthcare. However, most of the current research lacks input from end users, and therefore their practical value is limited. To address this, we conducted semi-structured interviews with clinicians to discuss their thoughts, hopes, and concerns. Clinicians from our sample generally think positively about developing AI-based tools for clinical practice, but they have concerns about how these will fit into their workflow and how it will impact clinician-patient relations. We further identify training of clinicians on AI as a crucial factor for the success of AI in healthcare and highlight aspects clinicians are looking for in (X)AI-based tools. In contrast to other studies, we take on a holistic and exploratory perspective to identify general requirements for (X)AI products for healthcare before moving on to testing specific tools.
@@ 1 Introduction
Rapid developments in artificial intelligence (AI) call for more transparency of AI-based technologies to tackle concerns related to trustworthiness, responsibility, and accountability. This has led to the emergence of the field of Explainable AI (XAI). The goal of XAI is making AI models more transparent and understandable to address these concerns and aid the implementation of AI-based technologies. Especially in high-stakes situations, like those in healthcare, AI-based systems could enhance critical decision-making processes (e.g., Eke and Shuib, 2024; Alowais et al., 2023). Ideally, AI models provide a certain prediction (e.g., the probability of a skin lesion being cancerous) which supports the clinicians in their decision process. However, implementation in the medical field – beyond regulatory requirements these technologies have to meet – has been slow partly due to lacking transparency (Eke and Shuib, 2024; He et al., 2019). Although user studies, evaluations, and multidisciplinary research are generally advocated for within XAI, only a small fraction of publications evaluate their proposed solutions with users. For example, Saarela and Podgorelec (2024) reviewed 512 articles, where almost 60% did not do any evaluation at all, and only 8% performed qualitative evaluation involving domain experts. Similarly, Nauta et al. (2023) report that only one in five papers, of over 300 papers they reviewed, performed evaluations with users. Arguably, XAI approaches to medical imaging as used in radiology may be most frequently implemented, as shown by the number of publications with XAI applications in this sub-field alone (e.g., Borys et al., 2023; Gupta and Seeja, 2024). For tabular data, however, there are less XAI applications, implementations and user tests. Therefore, our work focuses on tabular data structures like electronic health records. Specifically, in this research, we conduct semi-structured interviews with clinicians from the Netherlands. Our aim is to elicit clinicians’ thoughts on XAI, and by that, uncover compelling perspectives that may give rise to new research avenues and considerations in the design of XAI methods for healthcare and especially for tabular data structures. For example, feature importance scores are often praised for their simplicity, however our research suggests that they may become unbelievable due to exactly this property. Furthermore, our interviewees also raised concerns about taking responsibility for decisions, even if there is sufficient trust in the AI.
Overall, our interview study provides compelling statements and unique angles that put existing research into a new light and may give rise to new research avenues. We also find support for previous research, including that users emphasize the importance of receiving information about the background of the model in addition to its predictions and explanations. Our results further highlight the importance of multidisciplinary projects to ensure seamless introduction of (X)AI-based tools into clinical practice, and the significance of training clinicians on AI before implementing AI-based tools in their workflow.
@@ 2 Background
One can generally distinguish between glass-box and black-box models in XAI. Glass-box models are transparent, and model internals and operations are accessible to the users so that they can trace back every prediction made by the model. Thus, glass-box models provide both local and global explainability. Black-box models are opaque machine learning models where the user is unable to trace back the prediction due to the model complexity. Typically, (only) local explainability is achieved by complementing single predictions made by black-box models with post-hoc explanations. For an in-depth account of developments in XAI, we refer to several recent review papers (see, for example, Das et al., 2020; Gilpin et al., 2018; Murdoch et al., 2019; Ras et al., 2022; Linardatos et al., 2021; Guidotti et al., 2018; Carvalho et al., 2019; Barredo Arrieta et al., 2020; Ali et al., 2023)
@@
Healthcare is a high-stakes domain with vast opportunities for successful (X)AI implementations. This is also reflected in the increasing number of publications on AI-based clinical decision support tools (DSTs) that incorporate explainability. For example, Thoral et al. (2021) trained an AI-model to predict readmission or death within seven days after discharge from the ICU and explain predictions using feature importance scores (Lundberg and Lee, 2017). Both Abdullah et al. (2021) and Antoniadi et al. (2021) provide reviews of explainable machine learning approaches in healthcare including examples of publications on medical AI tools with explainability. Markus et al. (2021) focus on how explainability can be evaluated and provide a step-by-step guide to choose the most appropriate XAI method. Amann et al. (2020) take a multidisciplinary approach to determine the need for explainability in healthcare from technological, legal, medical, patient, and ethical perspectives. They conclude that there is an apparent need for explainability and highlight the importance of the interaction between healthcare professionals and AI-based DSTs.
@@
Cai et al. (2019) conduct a user study including semi-structured interviews on AI-assisted diagnosis of prostate cancer, focusing on identifying desired aspects for explainability in the AI tool. They find that pathologists are interested in global properties of the AI tool beyond accuracy, such as how conservative or liberal its decisions are or what kind of data it was trained on. In a similar fashion, Naiseh et al. (2023) empirically evaluate the effect of different types of explanations on trust calibration in a medical context and also conduct interviews with a smaller group of participants. Like Cai et al. (2019), they find that practitioners are interested in global information of AI-tools, such as the contents of the data and the recency of the trained model. They further find that practitioners worry about fitting AI-tools into their workflow because it adds extra workload and that they would like more guidance on how to interpret the explanations. Tonekaboni et al. (2019) conduct interviews with 10 clinicians that are familiar with AI-based clinical tools, primarily to identify which explanation types best serve their needs and to propose evaluation metrics of explainability tools. For example, they find that clinicians value feature importance scores as they draw attention to specific patient characteristics relevant for the prediction. They further conclude that example-based explanations are less helpful as they depend on the definition of similarity with the added difficulty that seemingly similar patients may have significantly different outcomes and vice versa. They also find that accuracy alone is not sufficient and that the output should be complemented with a certainty score for the prediction. Other qualitative studies address, for example, onboarding material for XAI-tools (Lee et al., 2024), adopt a broader stakeholder perspective (Subramanian et al., 2024), or analyze collaboration between developers and clinicians (Bienefeld et al., 2023).
@@
Our study differs from this previous research in several ways. First, we do not build our work around a specific case or AI-tool. In doing so, we eliminate the factor of expertise with the specific device or case from the study design. Besides, while we also consider this type of research important, we think it is just as relevant to take a step back and use a holistic approach for XAI in healthcare. We aim to elicit and analyze perspectives of clinicians that may be universal to a range of AI products and fields within healthcare. Second, in line with that, we interview clinicians from different fields within healthcare to record general data. Third, our study is exploratory in nature rather than empirically driven. That is, we do not test for an effect of XAI methods on, for example, trust or understanding. Qualitative interview studies are known to be exploratory in nature and generalizing findings may be difficult. Nonetheless, they pose a useful tool to gather in-depth understanding that can enrich results of quantitative studies and to uncover unique perspectives and compelling angles for further research (Tenny et al., 2022; Dunwoodie et al., 2023). The goal of our research is to do just that, as well as to identify clinicians’ requirements for successful implementation of AI-tools in clinical practice with the aim of improving decision making in healthcare.
@@ 3 Methodology
We conducted 11 semi-structured interviews with open-ended questions with clinicians trained and working in the Netherlands. In this section we discuss recruitment strategy, interview procedure, materials, and limitations of our design. The supplementary materials for the interviews are available in the appendix. The study was approved by the ethical body of our institution [the details will be filled after the review process in order to preserve the anonymity of our submission]. Interviewees were informed of the goal and procedure of the interviews and signed an informed consent form prior to their participation.
3.1 Participants and recruitment
The main inclusion criterion for participants was successful completion of their medical study in the Netherlands. The sample included clinicians from different fields within medicine (general practice, (pediatric) oncology, radiology, general health psychology, obstetrics and gynecology, surgery) and with years of experience in their fields ranging from two to 30 years. Participants were recruited through networking and word of mouth. None of the interviewees was personally known to the interviewer before the interview was conducted.
3.2 Procedure
Potential interviewees were contacted via email, which stated the reason for reaching out, the person that had provided the contact details, practical information such as the expected time, and the informed consent form. If interviewees responded positively, an exchange followed to arrange a meeting. Interviews were scheduled for 30 minutes, and the majority of them took place online as participants come from across the Netherlands. During the interview, the interviewer confirmed with participants that they agree to have the interview audio-recorded. Interviewees could ask questions or withdraw from the interview at any moment, and were explicitly given the opportunity to add any information they wished to mention or ask questions at the end of the interview. All interviews were conducted by the same interviewer to avoid potential differences due to interviewing style.
3.3 Materials
An interview scheme with guiding questions was developed to provide structure and guidance to the interviews (see Table 1). Broadly, aspects that we looked at include (i) tasks that clinicians (would) find (X)AI helpful for, (ii) desirable characteristics of XAI methods/explanations, (iii) hopes as well as concerns and fears they face regarding the implementation of (X)AI in their workflow, and (iv) their perception of the role of trust in the implementation and adoption of (X)AI.
@@ 3.3 Materials
At the start of each interview, the interviewer provided context for the interview after introductions, similar to what participants had read in the informed consent form beforehand. Then, the first set of questions was aimed at inquiring about the interviewee’s background, including their specialization and experience. The following questions were aimed at understanding the current use of AI in their field and healthcare in general. Besides, we also asked interviewees to share about research in (X)AI for healthcare that they are aware of. Then, we encouraged them to come up with use cases of AI in their field which they would find helpful in their job. Following up on that, we asked them for reasons why these do not yet exist, and why implementation of AI in healthcare in general is slow. The next set of questions was focused on uncovering interviewee’s attitude towards AI in healthcare. This entailed asking specifically for their positive and negative feelings towards using AI in healthcare, and if they could describe their hopes and concerns. In this part of the interview, we also asked them about the role they believe trust plays in the implementation and adoption of (X)AI in practice. Then, for the next set of questions, the interviewer explained three different types of explanations using a simple illustrative example for illustration (see Appendix A for the exact material and information given to participants) and asked interviewees to explain their thoughts considering each of these explanation types. The explanations that we chose were (1) feature importance scores, (2) counterfactual explanations, and (3) a glass box model.
Feature importance scores.
Feature importance methods calculate a score for each input feature that shows its contribution to the predicted score. We decided to include this method as it is arguably the most popular method within XAI at this moment. For our illustrative example, we utilized scores and the respective bar plot produced by SHAP (Lundberg and Lee, 2017). Starting from the mean prediction in the data, a positive SHAP value increases the score while a negative value decreases the score. Due to the additive nature of SHAP, adding up the SHAP values of all features is equivalent to the final prediction.
Counterfactual explanations.
Counterfactual explanations show the minimal change required in the input feature space to change the prediction of a sample to, for example, a different class. In the literature, the sample and its corresponding prediction that is to be explained is coined factual, whereas the changed data point leading to the desired prediction is called counterfactual. Research from the social and cognitive sciences shows that counterfactual explanations mimic the way humans give explanations themselves, and hence are expected to be good explanations in XAI (Byrne, 2019; Miller, 2019), which has spurred many researchers to propose algorithms generating these explanations (e.g., Russell, 2019; Ustun et al., 2019; Kanamori et al., 2020, 2021; Karimi et al., 2020, 2021; Mothilal et al., 2020; Mahajan et al., 2019; Maragno et al., 2022).
Rules as glass-box model.
As a contrast to the previous two explanation methods for black-box models, we opted for a glass-box model where the model itself is transparent and predictions can be traced back. For our illustrative example, we chose the method proposed by Röber et al. (2024), which produces a set of short and weighted rules. Each sample is classified by a subset of the rules, and rules receive weights which can be interpreted as importance. The final prediction stems from a weighted combination of the rule predictions.
Interviewees received a concise but detailed information sheet as supplementary material before the interview, which included the illustrative example and the three explanations, complemented with information about what they show and how to interpret them. We include the material that interviewees received in Appendix A. Besides their opinion on these specific explanation types, we included a question on what characteristics (e.g., visual, interactive) they would like to see in an AI system, or what type of information they would like to be shown. Then, at the end of the interview, interviewees were asked if they would like to share any other considerations with us.
3.4 Analysis
Upon completion of all interviews, each recording was transcribed and subsequently analyzed using inductive thematic analysis (Braun and Clarke, 2006). Inductive thematic analysis is a flexible methodology for analyzing qualitative data in a bottom-up manner, i.e., not driven by existing theory. In this data-driven approach, interviews are coded systematically by identifying features relevant for each aspect of the interview scheme and the overall research goal, collated by data relevant to each code. This step helps to compress the data into digestible chunks that at the same time retain the relevant information. Subsequently, themes are identified by grouping codes together, by that forming broader patterns of meaning related to the research goal. In this bottom-up and exploratory analysis, no predefined coding categories or themes are used; instead, responses are carefully analyzed to identify recurring themes and patterns and the codes are developed organically in response to the data.
3.5 Limitations
The nature of our study bears some limitations we should consider when interpreting the results. First, interviews are known to bear threats to validity and reliability (Alshenqeeti, 2014). As a response to this, we chose to have the same interviewer conduct all interviews. Further, interviews were audio-recorded and then transcribed for the analysis, by that avoiding misinformation through note-taking or incorrect understanding by the interviewer. Furthermore, during the interviews, time was taken to clarify any unclarities to avoid misunderstanding and misinterpretation of the responses. Second, as an explorative interview study, the sample size is limited and we did not perform any quantitative analyses of the results and should be careful with making generalizations. Third, the interviews were scheduled to last 30 minutes, which may not be considered long enough to discuss different aspects of XAI. As this is an exploratory study and clinicians are known to have limited time, we were constrained by this duration. Lastly, we reached out to candidates via email, and in some cases a common connection made the introduction. Generally, it can be expected that individuals interested in XAI are more likely to positively respond to an interview request. Hence, we may only have interviewed individuals with strong (positive or negative) opinions on the topic, by that missing out on the broader group of clinicians.
@@ 4 Results
In this section we discuss patterns identified through our thematic analysis of the interviews and put them into perspective using direct quotes and examples from interviewees.
@@ 4.1 Current use of AI in healthcare
Interviewees generally talked about AI-based tools in healthcare in three different categories, namely tools related to (1) business operations, (2) care processes, and (3) medical device regulations. The first category of AI systems is relatively universal to any larger business, not only hospitals or care facilities, and covers aspects like finance, forecasting, and so on. The second category relates to AI systems that enhance the care process by, for instance, making it more efficient. A typical example that several interviewees provided was the planning of chemotherapy. The last category of AI systems in healthcare concerns AI systems that fall under the medical device regulation. Simply put, a medical device is any instrument intended to be used to diagnose, treat, monitor, or prevent diseases “and does not achieve its primary intended action by pharmacological, immunological, or metabolic means” (International Medical Device Regulators Forum (2013), IMDRF, p.7). AI models fall under the category Software as a Medical Device, which includes software that is intended to be used for any of the aforementioned purposes without being part of a hardware medical device (International Medical Device Regulators Forum (2013), IMDRF). In layman’s terms, decision support tool is more commonly used to describe AI as a medical device. Here, an AI model is trained to make a prediction based on patient data and hence, supports clinicians in their decisions.
When asked to what extent AI models are implemented in medical care in their field specifically, interviewees unanimously said that there is very little AI (as DSTs) integrated in their workflows. Most interviewees pointed out that a significant amount of budget is invested in research and development of medical AI models. Although many of these efforts are being tested, very few of them actually reach the bedside. Arguably, the most popular field for AI applications within healthcare is radiology, with the number of publications on AI in radiology being “sky high” (interview 4). However, even though there are more than 200 products in AI for radiology available on the European market [^1] https://radiology.healthairegister.com, not many of those are integrated in clinical practice yet. Furthermore, one interviewee pointed out that most of these products perform specific subtasks, such as segmentation or quantification tasks, instead of predicting a concrete medical outcome; thus, only indirectly supporting clinicians.
@@ 4.2 What tasks should AI be used for?
Regarding AI as a clinical DST, four interviewees mentioned predicting complications as a useful application. For example, when performing surgery, they would like to know upfront how large the chance is to run into a certain complication. And even beyond that, one interviewee said they “would like to know upfront what can go wrong and how can we prevent that” (interview 5). Hence, AI could serve not only as a warning system but also as a guide on how to navigate certain situations. To give a concrete example, one interviewee from obstetrics and gynecology explained that when a woman is in labor, they monitor both the mother and the baby, meaning they collect a large dataset. Having an AI model trained on this data that continuously predicts the chance of a birth ending in a C-section would save them not only time and resources, but would also reduce the risks for women in labor. Another example comes from oncology, where it would be beneficial to predict the likelihood of complications during surgery but also the probability of remission following treatment.